rented 2 movies and
5
video games for a total of
$35
. The next month he rented
8
movies and
3
video games for a total of
$38
. Find the rental cost for each movie and each video game.

Answer by Cromlix(4381)

Hi there,
2 movies + 5 games = 35......(1)
8 movies + 3 games = 38......(2)
Multiply (1) by 4
8 movies + 20 games = 140....(1)
8 movies + 3 games = 38......(2)
Subtract (2) from (1)
..........17 games = 102
...........1 game = $6
Substitute game = $6 into (2)
8 movies + 3 games = 38......(2)
8 movies + 3(6) = 38
8 movies + 18 = 38
8 movies = 38 - 18
8 movies = 20
1 movie = 20/8
1 movie = $2.50
Hope this helps :-)